OpenStudy (magbak):

PLEASE HELP ME!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! I WILL AWARD MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Given the sequence 3, 9, 15, which of the following is equivalent to (look at equation posted below) Answer Choices 54 102 144 168

OpenStudy (anonymous):

wheres the equation?

OpenStudy (magbak):

\[\sum_{n=5}^{8}a_{n}\]

OpenStudy (magbak):

@primeralph

OpenStudy (primeralph):

@CarlosGP will help you.

OpenStudy (magbak):

come on it is always good to have a second opinion.

OpenStudy (anonymous):

\[a_n=3+6(n-1)=6n-3\] \[a_5=27\]\[a_8=45\] \[S=\frac{ 27+45 }{ 2 } \times 4=144\]

OpenStudy (magbak):

Is that right @primeralph

OpenStudy (anonymous):

@magbak if you dont trust my solution I can delete it

OpenStudy (primeralph):

@magbak Sometimes, understand that the other person may be smarter than me too.

OpenStudy (magbak):

no I trust it ok. I am sorry it is nothing personal just my parents taught me to do that @CarlosGP

OpenStudy (primeralph):

In my opinion, it's right.

OpenStudy (magbak):

I am so sorry.

OpenStudy (anonymous):

mine taught me to be confident

OpenStudy (magbak):

Me too. I am so sorry.

OpenStudy (anonymous):

never mind